The sum of Joan's age and her fathers age is 67. Three times Joan's age increased by 7 is her fathers age. How old is Joan?
Answer by addingup(3677)

J + f = 67 subtract J from both sides:
f = 67 - J
--------------------------
3J + 7 = f substitute for f:
3J + 7 = 67 - J
4J = 60
J = 15
and f?
f = 67 - J = 67 - 15 = 52
----------------------------------
Check:
f = 3J + 7
f = 3(15) + 7
f = 45 + 7 = 52 Correct